OpenStudy (anonymous):

Find the x-intercept of -4x^2-7x+4 and how?

OpenStudy (anonymous):

factor

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=1T-rsltsWnM

OpenStudy (anonymous):

x-intercepts means (x1,0) and (x2,0) -4x^2-7x+4=0  plug in values of a,b and c x1=(1/8)(-7-sqrt(113)) x2=(1/8)(sqrt(113)-7)